# Kernel-U-Net: Hierarchical and Symmetrical Framework for Multivariate Time Series Forecasting

### 03 Jan 2024 v1

@@ 3.2 Kernel U-Net
1:  # Create a list of multiples
@@
3:  # Create first layer
4:  layers = [KernelWrapper($f_{n+m},$ $1,$ $M_{h},$ $L_{n+m},$ $M_{n+m}$)]
5:  # Create intermediate layers
6:  for  $i$ in $[n+m-1,...,n+2,n,...,2]$ do
[…]
8:  end for
9:  # Create last layer
10:  layers.append(KernelWrapper($f_{2},1,$ $M_{2},$ $L_{1},$ $M_{1})$)
11:  decoder = nn.Sequential (*layers)
Kernel Wrapper As described previously, Kernel U-Net accepts custom kernels that follow predefined design patterns. A kernel wrapper is an interface that executes the given kernel inside the encoder and decoder. It requires necessary parameters such as model kernel, input dimension, input length, output dimension, and output length and holds global variables such as skip connection vector. By default, a kernel reduces the length of a segment from $L_{i}$ to $1$ in the encoder and develops the length of a segment from $1$ to $L_{i}$ in the decoder. The kernel wrapper initiates an instance of a given kernel and calls it in the forward function.
@@
1:  # Initiation of encoder and decoder in K-U-N.
2:  encoder = KernelUNetEncoder(inputs)
3:  decoder = KernelUNetDecoder(inputs)
4:
5:  # Define the forward function as
6:  z = encoder(x)
7:  for  $i$ in $[1,...,n]$ do
8:     output = encoder.layers[i]._unet_skip_output
9:     decoder.layers[$n-i+1$]. _unet_skip_input = output
10:  end for
11:  y= decoder(z)
12:  return y
Creation of Kernel U-Net
Let us note input length $L$, input dimension $M$, Lists of multiples of look-back window $,[L_{2},$...,… , $L_{n}]$ and feature $,[L_{n+2},$...,… , $L_{n+m}]$ , size of segment-unit $L_{1}$ and feature-unit $M_{1}$ , a list of hidden dimension $[M_{2},...,M_{n},$ $M_{n+2},...,M_{n+m}]$, a list of kernels $[f_{1},f_{2},...,f_{n},$ $f_{n+2},...,f_{n+m}]$, latent vector length $L_{h}$ and dimension $M_{h}$.
Remark that the hidden dimension of the model is the size of intermediate output vectors within layers. It corresponds to channel size in a convolutional network. This dimension can be augmented for a larger passage of information. The kernel wrapper is described in algorithm 1. The creation of Kernel U-Net Encoder, Decoder, and complete U-Net are described in algorithms 2, 3, 4.
Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer Informer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
96 0.369 0.404 0.37 0.4 0.374 0.394 0.375 0.399 0.376 0.419 0.449 0.459 0.865 0.713
192 0.397 0.426 0.413 0.429 0.408 0.415 0.405 0.416 0.42 0.448 0.5 0.482 1.008 0.792
336 0.429 0.428 0.422 0.44 0.429 0.427 0.439 0.443 0.459 0.465 0.521 0.496 1.107 0.809
720 0.433 0.45 0.447 0.468 0.44 0.453 0.472 0.49 0.506 0.507 0.514 0.512 1.181 0.865
ETTh2
96 0.271 0.336 0.274 0.337 0.277 0.338 0.289 0.353 0.346 0.388 0.358 0.397 3.755 1.525
192 0.332 0.376 0.339 0.379 0.344 0.381 0.383 0.418 0.429 0.439 0.456 0.452 5.602 1.931
336 0.357 0.4 0.329 0.384 0.357 0.4 0.448 0.465 0.496 0.487 0.482 0.486 4.721 1.835
720 0.39 0.438 0.379 0.422 0.394 0.436 0.605 0.551 0.463 0.474 0.515 0.511 3.647 1.625
ETTm1
96 0.285 0.337 0.29 0.342 0.306 0.348 0.299 0.343 0.379 0.419 0.505 0.475 0.672 0.571
192 0.327 0.367 0.332 0.369 0.349 0.375 0.335 0.365 0.426 0.441 0.553 0.496 0.795 0.669
336 0.357 0.392 0.366 0.392 0.375 0.388 0.369 0.386 0.445 0.459 0.621 0.537 1.212 0.871
720 0.408 0.422 0.416 0.42 0.433 0.422 0.425 0.421 0.543 0.49 0.671 0.561 1.166 0.823
ETTm2
96 0.16 0.245 0.165 0.255 0.167 0.255 0.167 0.26 0.203 0.287 0.255 0.339 0.365 0.453
192 0.215 0.215 0.22 0.292 0.221 0.293 0.224 0.303 0.269 0.328 0.281 0.34 0.533 0.563
336 0.268 0.326 0.274 0.329 0.274 0.327 0.281 0.342 0.325 0.366 0.339 0.372 1.363 0.887
720 0.348 0.378 0.362 0.385 0.368 0.384 0.397 0.421 0.421 0.415 0.433 0.432 3.379 1.338
Electricity
96 0.129 0.225 0.129 0.222 0.141 0.237 0.14 0.237 0.193 0.308 0.201 0.317 0.274 0.368
192 0.148 0.244 0.147 0.24 0.154 0.248 0.153 0.249 0.201 0.315 0.222 0.334 0.296 0.386
336 0.168 0.263 0.163 0.259 0.171 0.265 0.169 0.267 0.214 0.329 0.231 0.338 0.3 0.394
720 0.205 0.299 0.197 0.29 0.21 0.297 0.203 0.301 0.246 0.355 0.254 0.361 0.373 0.439
Traffic
96 0.360 0.252 0.36 0.249 0.41 0.279 0.41 0.282 0.587 0.366 0.613 0.388 0.719 0.391
192 0.377 0.263 0.379 0.25 0.423 0.284 0.423 0.287 0.604 0.373 0.616 0.382 0.696 0.379
336 0.389 0.270 0.392 0.264 0.435 0.29 0.436 0.296 0.621 0.383 0.622 0.337 0.777 0.42
720 0.427 0.290 0.432 0.286 0.464 0.307 0.466 0.315 0.626 0.382 0.66 0.408 0.864 0.472
Weather
96 0.146 0.200 0.149 0.198 0.182 0.232 0.176 0.237 0.217 0.296 0.266 0.336 0.3 0.384
192 0.192 0.244 0.194 0.241 0.225 0.269 0.22 0.282 0.276 0.336 0.307 0.367 0.598 0.544
336 0.241 0.287 0.245 0.282 0.271 0.301 0.265 0.319 0.339 0.38 0.359 0.395 0.578 0.523
720 0.310 0.340 0.314 0.334 0.338 0.348 0.323 0.362 0.403 0.428 0.419 0.428 1.059 0.741
3.3 Custom Kernels
Linear kernel
The linear kernel is a simple matrix multiplication. Given $X$ of shape $(B,L_{i},M_{i})$, linear kernel $f_{i}$ reshape it into $(B,L_{i}*M_{i})$ and process it as follow:
$Z_{i}=f_{i}(X)=X*w+b$
, where $Z_{i}$ is output of kernel, $w\in R^{L_{i}M_{i}\times L_{i+1}M_{i+1}}$ is weight matrix and $b\in R^{L_{i+1}M_{i+1}}$ is bias vector. Remark that linear Kernel U-Net is an equivalent version to convolutional U-Net whose stride size is $L_{i}$. We also test a non-linear kernel that takes a non-linear activation function such as Tanh in the intermediate hidden layers.

### 12 Feb 2024 v2

@@ 3.2 Kernel U-Net
Kernel Wrapper As described previously, Kernel U-Net accepts custom kernels that follow predefined design patterns. A kernel wrapper inherits ”nn.Module” class in Pytorch and offers an interface that executes the given kernel inside the encoder and decoder. It requires necessary parameters such as model kernel, input dimension, input length, output dimension, and output length and holds global variables such as skip connection vector (Algorithm 1). By default, a kernel receives a batch of patches in the shape of $(-1,L_{i},M_{i})$ and outputs a new batch of patches in the shape of $(-1,L_{i+1},M_{i+1})$. Remark that the kernels reduce the length of patches from $L_{i}$ to $1$ in the encoder and expand the length of patches from $1$ to $L_{i}$ in the decoder. The kernel wrapper initiates an instance of a given kernel and calls it in the forward function. It reshapes the input patches and checks the output shape outside of the kernel.
Algorithm 2 Kernel U-Net Encoder
Input: $L_{1},M_{1}$, $[L_{2},$,...,… ,$L_{n},$ $L_{n+2},$,...,… ,$L_{n+m}]$, $[M_{2},$,...,… ,$M_{n},$ $M_{n+2},$,...,… ,$M_{n+m}]$, $[f_{1},f_{2},$,...,… ,$f_{n},$ $f_{n+2},...,f_{n+m}]$, $L_{h}$, $M_{h}$
Output: Instance of Kernel U-Net Encoder
  # Define the init function :
  def __init(inputs)__:
  $L=[L_{1},L_{2},...,L_{n},$ $L_{n+2},...,L_{n+m}]$
  # Create first layer
[…]
  # Create intermediate layers
[…]
           layers.append(KernelWrapper($f_{i},$ $L_{i},$ $M_{i},$ $1,$
           $M_{i+1}$))
  end for
  # Create last layer
  layers.append(KernelWrapper($f_{n+m},L_{n+m},$
  $M_{n+m},$ $L_{h},M_{h}$))
  encoder = nn.Sequential (*layers)
Creation of Kernel U-Net
Let us note input length $L$, input dimension $M$, Lists of multiples of look-back window $,[L_{2},$...,… , $L_{n}]$ and feature $,[L_{n+2},$...,… , $L_{n+m}]$ , size of patch $L_{1}$ and feature-unit $M_{1}$ , a list of hidden dimension $[M_{2},...,M_{n},$ $M_{n+2},...,M_{n+m}]$, a list of kernels $[f_{1},f_{2},...,f_{n},$ $f_{n+2},...,f_{n+m}]$, latent vector length $L_{h}$ and dimension $M_{h}$ Remark that the hidden dimension of the model is the size of intermediate output vectors within layers. It corresponds to channel size in a convolutional network. This dimension can be augmented for a larger passage of information. We describe the creation of the K-U-Net encoder in Algorithm 2 and leave that of the decoder (Algorithm 3) and K-U-Net (Algorithm 4) in the appendix.
3.3 Custom Kernels
Linear kernel
The linear kernel is a simple matrix multiplication (Figure 3). Given $X$ of shape $(B,L_{i},M_{i})$, linear kernel $f_{i}$ reshape it into $(B,L_{i}*M_{i})$ and process it as follow:
$Z_{i}=f_{i}(X)=X*w+b$
, where $Z_{i}$ is output of kernel, $w\in R^{L_{i}M_{i}\times L_{i+1}M_{i+1}}$ is weight matrix and $b\in R^{L_{i+1}M_{i+1}}$ is bias vector. Remark that the number of parameters of $w$ in a linear kernel is $L_{i}*1*ch_{in}*ch_{out}$ which is equivalent to that of a 1D convolutional layer whose kernel size is $L_{i}*1$.
The multi-layer linear kernel has an additional hidden layer and a non-linear activation function Tanh. The formulation is:
$Z_{i}=f_{i}(X)=Tanh(X*w_{1}+b_{1})*w_{2}+b_{2}$
, where $Z_{i}$ is output of kernel, $w_{1}\in R^{L_{i}M_{i}\times L^{\prime}M^{\prime}}$ and $w_{1}\in R^{L^{\prime}M^{\prime}\times L_{i+1}M_{i+1}}$ are weight matrix , $b_{1}\in R^{L^{\prime}M^{\prime}}$ and $b_{2}\in R^{L_{i+1}M_{i+1}}$ are bias vectors, and $L^{\prime}=\frac{1}{2}(L_{i}+L_{i+1}),M^{\prime}=\frac{1}{2}(M_{i}+M_{i+1})$.
Transformer kernel
The Vanilla Transformer is made of a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with Relu activation. The transformer kernel in this work follows the classical structure.
LSTM kernel
The LSTM kernel contains a classic LSTM cell and a linear layer for skip connection. The hidden states of all time steps are flattened and used for a linear combination in the next. The detailed structure of the Transformer kernel and the LSTM kernel is discussed in the appendix.
Figure 3: Structure of Linear Kernel. a) Linear kernel, b) Multi-layer Linear Kernel with Tanh activation.
Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer Informer Yformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
96 0.370 0.393 0.37 0.4 0.374 0.394 0.375 0.399 0.376 0.419 0.449 0.459 0.865 0.713 0.985 0.74
192 0.404 0.414 0.413 0.429 0.408 0.415 0.405 0.416 0.42 0.448 0.5 0.482 1.008 0.792 1.17 0.855
336 0.420 0.430 0.422 0.44 0.429 0.427 0.439 0.443 0.459 0.465 0.521 0.496 1.107 0.809 1.208 0.886
720 0.438 0.454 0.447 0.468 0.44 0.453 0.472 0.49 0.506 0.507 0.514 0.512 1.181 0.865 1.34 0.899
ETTh2
96 0.271 0.335 0.274 0.337 0.277 0.338 0.289 0.353 0.346 0.388 0.358 0.397 3.755 1.525 1.335 0.936
192 0.332 0.377 0.339 0.379 0.344 0.381 0.383 0.418 0.429 0.439 0.456 0.452 5.602 1.931 1.593 1.021
336 0.357 0.4 0.329 0.384 0.357 0.4 0.448 0.465 0.496 0.487 0.482 0.486 4.721 1.835 1.444 0.96
720 0.39 0.438 0.379 0.422 0.394 0.436 0.605 0.551 0.463 0.474 0.515 0.511 3.647 1.625 3.498 1.631
ETTm1
96 0.286 0.342 0.29 0.342 0.306 0.348 0.299 0.343 0.379 0.419 0.505 0.475 0.672 0.571 0.849 0.669
192 0.330 0.363 0.332 0.369 0.349 0.375 0.335 0.365 0.426 0.441 0.553 0.496 0.795 0.669 0.928 0.724
336 0.360 0.384 0.366 0.392 0.375 0.388 0.369 0.386 0.445 0.459 0.621 0.537 1.212 0.871 1.058 0.786
720 0.405 0.412 0.416 0.42 0.433 0.422 0.425 0.421 0.543 0.49 0.671 0.561 1.166 0.823 0.955 0.703
ETTm2
96 0.16 0.245 0.165 0.255 0.167 0.255 0.167 0.26 0.203 0.287 0.255 0.339 0.365 0.453 0.487 0.529
192 0.215 0.215 0.22 0.292 0.221 0.293 0.224 0.303 0.269 0.328 0.281 0.34 0.533 0.563 0.789 0.705
336 0.268 0.326 0.274 0.329 0.274 0.327 0.281 0.342 0.325 0.366 0.339 0.372 1.363 0.887 1.256 0.904
720 0.343 0.379 0.362 0.385 0.368 0.384 0.397 0.421 0.421 0.415 0.433 0.432 3.379 1.338 2.698 1.297
Electricity
96 0.129 0.226 0.129 0.222 0.141 0.237 0.14 0.237 0.193 0.308 0.201 0.317 0.274 0.368 - -
192 0.147 0.244 0.147 0.24 0.154 0.248 0.153 0.249 0.201 0.315 0.222 0.334 0.296 0.386 - -
336 0.163 0.261 0.163 0.259 0.171 0.265 0.169 0.267 0.214 0.329 0.231 0.338 0.3 0.394 - -
720 0.197 0.292 0.197 0.29 0.21 0.297 0.203 0.301 0.246 0.355 0.254 0.361 0.373 0.439 - -
Traffic
96 0.358 0.253 0.36 0.249 0.41 0.279 0.41 0.282 0.587 0.366 0.613 0.388 0.719 0.391 - -
192 0.373 0.262 0.379 0.25 0.423 0.284 0.423 0.287 0.604 0.373 0.616 0.382 0.696 0.379 - -
336 0.390 0.271 0.392 0.264 0.435 0.29 0.436 0.296 0.621 0.383 0.622 0.337 0.777 0.42 - -
720 0.430 0.292 0.432 0.286 0.464 0.307 0.466 0.315 0.626 0.382 0.66 0.408 0.864 0.472 - -
Weather
96 0.142 0.195 0.149 0.198 0.182 0.232 0.176 0.237 0.217 0.296 0.266 0.336 0.3 0.384 - -
192 0.188 0.244 0.194 0.241 0.225 0.269 0.22 0.282 0.276 0.336 0.307 0.367 0.598 0.544 - -
336 0.241 0.284 0.245 0.282 0.271 0.301 0.265 0.319 0.339 0.38 0.359 0.395 0.578 0.523 - -
720 0.310 0.333 0.314 0.334 0.338 0.348 0.323 0.362 0.403 0.428 0.419 0.428 1.059 0.741 - -
3.4 Complexity Analysis
As the K-U-Net is symmetric, we study the complexity of the encoder. Let us suppose that a kernel is receiving a sequence of length $L$ where $L=\prod_{1}^{n}L_{i}$ and all $L_{i}$ are equal. As the patch size in the first layer of Kernel U-Net encoder is $L_{1}$, a kernel module will process $\frac{L}{L_{1}}$ patches of size $L_{1}$. Therefore, the complexity is $O(\frac{L}{L_{1}^{n}}*f(L_{1}))$ where $f(L_{1})$ is the complexity inside the kernel in the function of the patch size $L_{1}$ and $n$ is the index of the layer. In the case of using the linear kernel at the first layer, the complexity is $O(\frac{L}{L_{1}}*L_{1})$= $O(L)$. In the case of using a classic transformer kernel, the complexity is $O(\frac{L}{L_{1}}*L_{1}^{2})$= $O(L*L_{1})$. Let us set $L_{1}=log(L)$, the complexity of the application of such a quadratic calculation kernel is $O(L*log(L))$. Moreover, if we apply the transformer kernel starting from the second layer, the complexity is dramatically reduced to $O(\frac{L}{L_{1}^{2}}*L_{1}^{2})$= $O(L)$. Following the same demonstration, the complexity of using multi-layer linear and LSTM kernels starting from the second layer is also bounded by $O(L)$.
Figure 4: Search result of K-U-Net with Multi-layer Linear, LSTM, Transformer kernels on ETTh1 dataset.
4 Experiments and Results
Datasets.
We conducted experiments on our proposed Kernel U-Net on 7 public datasets, including 4 ETT (Electricity Transformer Temperature) datasets [^1] https://github.com/zhouhaoyi/ETDataset (ETTh1, ETTh2, ETTm1, ETTm2), Weather [^2] https://www.bgc-jena.mpg.de/wetter/, Traffic [^3] http://pems.dot.ca.gov and Electricity [^4] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams
20112014
. These datasets have been benchmarked and publicly available onWu et al. (2021) and the description of the dataset is available in Zeng et al. (2023)).
Figure 5: Search result of K-U-Net with Multi-layer Linear, LSTM, Transformer kernels on Weather dataset.
Here, we followed the experiment setting in Zeng et al. (2023) and partitioned the data into $[12,4,4]$ months for training, validation, and testing respectively for the ETT dataset. The data is split into $[0.7,0.1,0.2]$ for training, validation, and testing for the Weather, Traffic, and Electricity datasets.
Baselines and Experimental Settings.
We follow the experiment setting in NLinear Zeng et al. (2023) and take $L=\{336,720\}$ step historical data as input then forecast $T\in\{96,192,336,720\}$ step value in the future. We replace the last value normalization with mean value normalization for ETT, Electricity, and Weather datasets, and apply instance normalizationUlyanov et al. (2017) for the Traffic dataset. We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) as mentioned in Wu et al. (2021).
We include recent methods: PatchTSTNie et al. (2023), NLinear, DLinearZeng et al. (2023), FEDformer Zhou et al. (2022), Autoformer Wu et al. (2021), Informer Zhou et al. (2021), LogTrans Li et al. (2019) YformerMadhusudhanan et al. (2023). We merge the result reported in Nie et al. (2023) and Zeng et al. (2023) for taking their best in a supervised setting and execute Yformer in our environment with default parameters in its Github [^5] https://github.com/18kiran12/Yformer-Time-Series-Forecasting.
Figure 6: Computation time of K-U-Net, PatchTST and Yformer.
Model Variants Search.
We propose 4 kernels for experiments with K-U-Net on 7 datasets. By replacing a linear kernel at different layers with other types of kernels for searching for variants that adapt the dataset. The variants are noted as ”K-U-Net $<$kernel$>$_$<$replacing_layer$>$ ($<$lookback_window$>$)”. For example, K-U-Net Linear_$\_0000(720)$ means that the model is made of linear kernels at all layers and the lookback window size is 720, K-U-Net Transf_$\_0100(720)$ means that a transformer kernel replaces the linear kernel at the second layer, K-U-Net Linear_$\_0110(720)$ means that multilayer linear kernels replace the kernels at the second and third layers. We compose 16 variants with Multi-layer linear kernels and 7 variants with Transformer and LSTM kernels.
To enumerate all variants that achieve at least once the best result, we report their performance by the average of the top 5 minimum running MSE(Top5MMSE) values on the validation and test set. The search results are reported with a relative score(RS) to the minimum Top5-MMSE:
$RS(m)=Min_{t\in T}\frac{Top5MMSE(m,t)}{Min_{\{m\in M,l\in L\}}Top5MMSE(m,t,l)}$
, where $m$ is a model in the models set $M$, $t$ is a forecasting horizon in $T$, $l$ is a look back window in $L$. Relative score notes the best-performed model with 1 and thus helps to identify the high-potential candidates for further fine-tuning examinations.
We observe in Figure 4 that the best variant for ETTh1 dataset is K-U-Net Linear_$\_0010$. Comparing the relative score of the K-U-Net of replacing__\__layer code ($0010$, $0011$), ($0110$, $0111$) and ($0100$, $0101$), we remark that replacing the highest layer with Transformer and LSTM kernel degrades the performance because of overfitting. Furthermore, We observe in Figure 5 that the best variants for the Weather dataset are K-U-Net Linear_$\_0110$, K-U-Net LSTM_$\_0100$ and K-U-Net Transf_$\_0100$. Comparing the relative score of the K-U-Net of replacing__\__layer code ($0010$, $0110$), ($0011$, $0111$) and ($0001$, $0101$) we remark that replacing the second layer with Transformer and LSTM kernel gains the performance for their expressiveness.
Among candidates in the search phase, we empirically choose 3 variants for fine-tuning experiments with 5 runs. The final result shows that the performance of Kernel U-Net exceeds or meets the state-of-the-art methods in univariate settings and multivariate settings in most cases.
Multivariate time series forecasting result.
We remark that our model improves the MSE performance around $72.92\%$ compared with Yformer and $0.99\%$ compared to PatchTST and NLinear in the multivariate setting (Table 1). It is worth noting that K-U-Net achieves similar results on the Electricity dataset with a variant based on Multilayer linear kernels.
Univariate time series forecasting result.
Additionally, we remark that our model improves the MSE performance around $60.47\%$ compared with Yformer and $6.06\%$ compared to PatchTST and NLinear in the Univariate setting(Table 3). It is worth noting that K-U-Net improves the MSE performance on ETTh2 for $15.67\%$ compared to PatchTST.
Figure 7: GPU consumation of K-U-Net, PatchTST and Yformer.
Computation Efficiency
We examined the computation efficiency of 4 variants of K-U-Net, PatchTST, and Yformer. We execute the models on the ETTh1 and ETTm1 datasets for 10 epochs and measure the average execution time per epoch and the GPU consumption during the training. The hidden dimension is set to 128 for all models. For fair comparison, PatchTST and Yformer contain 2 layers of Transformer block which equals to K-U-Net_Transf_$\_0100$. All experiments are executed on a Tesla V100 GPU in a Google Colab environment. Comparing to the PatchTST, K-U-Net Linear_$\_0000$ saves $85.20\%$ and $83.07\%$ memory (Figure 7) and saves $81.82\%$ and $80.55\%$ computation time (Figure 6) on ETTh1 and ETTm1 datasets, K-U-Net_Transf_$\_0100$ saves $55.15\%$ and $50.00\%$ memory and saves $28.47\%$ and $26.88\%$ computation time respectively.
5 Conclusion
In this paper, we propose Kernel-U-Net, a highly potential candidate for large-scale time series forecasting tasks. It provides convenience for composing particular models with custom kernels, thereby it adapts well to particular datasets. As an efficient architecture, it accelerates the procedure of searching for appropriate variants. Kernel-U-Net either exceeds or meets the state-of-the-art results in most cases. In the future, we hope to develop more kernels and hope that Kernel U-Net can be useful for other time series tasks such as classification or anomaly detection.
Appendix A Appendix
A.1 Complement Algorithms
The complete algorithm of Kernel U-Net contains creation of Encoder, Decoder (Algorithm 3), Wrapper, and Kernel U-Net(Algorithm 4).
@@ A.1 Complement Algorithms
  # Define the init function :
  def __init(inputs)__:
  # Create a list of multiples
@@
  # Create first layer
  layers = [KernelWrapper($f_{n+m},$ $1,$ $M_{h},$
  $L_{n+m},$ $M_{n+m}$)]
  # Create intermediate layers
  for  $i$ in $[n+m-1,...,n+2,n,...,2]$ do
           layers.append(KernelWrapper($f_{i},$ $1,M_{i+1}$
           $,L_{i},M_{i})$)
  end for
  # Create last layer
  layers.append(KernelWrapper($f_{2},1,$ $M_{2},$ $L_{1},$ $M_{1})$)
  decoder = nn.Sequential (*layers)
@@
  # Define the init function :
  def __init(inputs)__:
  # Initiation of encoder and decoder in K-U-N.
  encoder = KernelUNetEncoder(inputs)
  decoder = KernelUNetDecoder(inputs)
  # Define the forward function :
  def forward(x):
  reshape x to $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$
  transpose x to $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$
  reshape x to $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$
  # Apply encoder on x
  z = encoder(x)
  for  $i$ in $[1,...,n]$ do
           output = encoder.layers[i]._unet_skip_output
           decoder.layers[$n-i+1$]._unet_skip_input
           = output
  end for
  # Apply decoder on z
  y= decoder(z)
  reshape y to $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$
  transpose y to $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$
  reshape y to $(B,L,M)$
  return y
A.2 LSTM and Transformer kernels